# MACE-Dance: Motion-Appearance Cascaded Experts for Music-Driven Dance Video Generation

### v1

@@ 3.2 Motion Expert
Diffusion. DDPM [10] defines diffusion as a Markov noising process with latents $\{z_{t}\}_{t=0}^{T}$ that follow a forward noising process $q(z_{t}|x)$, where $x\sim p(x)$ is drawn from the 3D dance data distribution. The forward noising process is defined as:
$q(z_{t}|x)\sim\mathcal{N}(\sqrt{\bar{\alpha}_{t}}x,(1-\bar{\alpha}_{t})I),$ (1)
@@
where $\bar{\alpha}_{t}\in(0,1)$ are constants which follow a monotonically decreasing schedule such that when $\bar{\alpha}_{t}$ approaches 0. Timestep $T$ are commonly set to 1000, and $z_{T}\sim\mathcal{N}(0,I)$. With paired music conditioning $c$, we can reverse the forward diffusion process by learning to estimate $\hat{x}_{\theta}(z_{t},t,c)\approx x$ with model parameters $\theta$ for all $t$. We can optimize $\theta$ by the naive reconstruction loss in Diffusion Model [10]:
$\mathcal{L}_{\text{DM}}=\mathbb{E}\big[\|\hat{x}_{\theta}(z_{t},t,c)-x\|_{2}^{2}\big].$ (2)
Guidance-Free Training. While conventional Classifier-Free Guidance (CFG [9]) provides a knob for trading off diversity against fidelity, it is notoriously sensitive to the guidance scale. Suboptimal settings can easily disrupt this balance, manifesting in 3D motion as excessive jitter or repetitive, monotonous movements, ultimately degrading the dance quality. Moreover, the search for a reliable scale also incurs substantial tuning overhead. Motivated by these limitations, we adopt Guidance-Free Training (GFT [3]). GFT reformulates conditional training to directly learn a $\beta$-indexed sampling model via linear interpolation with the unconditional output. This allows a single model to capture an entire family of diversity-fidelity trade-offs robustly, eliminating the need for post-hoc guidance. Accordingly, we establish $x_{\beta}$ as the new optimization target for our model $\theta$:
$x_{\beta}=\beta\hat{x}_{\theta}(z_{t},t,c,\beta)+(1-\beta)\mathbf{sg}[\hat{x}_{\theta}(z_{t},t,\emptyset,1)]$ (3)
@@
$\mathcal{L}_{\text{rec}}$ $=\mathbb{E}\big[\|x_{\beta}-x\|_{2}^{2}\big],$ (4)
@@
$\mathcal{L}=\lambda_{\text{rec}}\mathcal{L}_{\text{rec}}+\lambda_{\text{joint}}\mathcal{L}_{\text{joint}}+\lambda_{\text{vel}}\mathcal{L}_{\text{vel}}+\lambda_{\text{foot}}\mathcal{L}_{\text{foot}}.$ (5)
@@
Inference. At each of the denoising timesteps $t$, Motion Expert predicts the denoised sample and noises it back to timestep t1t-1: $\hat{z}_{t-1}\sim q(\hat{x}_{\theta}(z_{t},t,c,\beta),t-1)$, terminating when it reaches $t=0$. We utilize Denoising Diffusion Implicit Models (DDIM [33]) to accelerate the sampling procedure. Values of $\beta$ near 0 favor high fidelity, while values near 1 favor high diversity. Thus, $\beta$ can also be regarded as a control signal, and we set its value to 0.75. Notably, GFT inherently achieves nearly double the generation efficiency compared to conventional CFG, as it only requires a single conditional computation per step, eliminating the need for simultaneous conditional and unconditional predictions.
3.2.2 Model Architecture
@@
Overview. Motion Expert adopts a BiMamba–Transformer hybrid backbone, thereby enabling the generation of temporally coherent and musically aligned dance motions. BiMamba captures intra-modal local dependencies in music or dance, while the Transformer models cross-modal global context. As shown in Fig.  2, the architecture details are as follows: Firstly, our model conditions the generator on the Librosa [22]-extracted music features from $M$ as [16], which are then processed by an $L_{m}$‑layer BiMamba to capture intra‑modal temporal dynamics. Secondly, the diffusion time step $t$ and temperature parameter $\beta$ are encoded as sinusoidal embeddings and fused by element-wise addition to yield a $t$-$\beta$ embedding used throughout the generator. Third, the dance generator consists of $L_{d}$ stacked blocks. In each block: (1) the current state $z_{t}$ is first passed through a BiMamba to model intra-modal local dependencies; (2) FiLM  [29] is applied to modulate the features with the fused $t$-$\beta$ embedding; (3) a Transformer performs cross-modal attention over the music encoding to integrate global musical context, and subsequently passes the result through a feed-forward network; and (4) a second FiLM [29] further reinforces the $t$-$\beta$ conditioning. Finally, the generator outputs the 3D motion sequence $\hat{x}_{\theta}(z_{t},t,c,\beta)$ (i.e. $X$), represented as SMPL [21] parameters. Owing to this architecture, the Motion Expert generates the entire sequence in a non-autoregressive manner during inference, not only improving generation efficiency but also avoiding the exposure-bias problem in autoregressive [47] and inpainting-based [37] methods.
@@
Intra-Modal Local-Dependency. While the Transformer excels at temporal modeling, it is inherently position-invariant and captures sequence order only through positional encodings [39], which limits its deep understanding of local dependencies. In contrast, music-to-dance generation demands strong local continuity between movements. Owing to its inherent sequential inductive bias, Mamba [8] has demonstrated strong performance in modeling fine-grained local dependencies  [42, 7]. Moreover, Bidirectional Mamba processes inputs in both forward and backward directions, enabling wider representations and deeper understanding of music and dance. Specifically, the Selective State Space Model (Mamba) integrates a selection mechanism and a scan module (S6) [8] to dynamically emphasize salient input segments for efficient sequence modeling. Unlike traditional SSMs with time-invariant parameters, Mamba generates input-dependent $\bar{A}_{t},\bar{B}_{t},C_{t}$ through fully connected layers, enhancing generalization. For each time step $t$, the input $x_{t}$, hidden state $h_{t}$, and output $y_{t}$ evolve as:
$h_{t}=\bar{A}_{t}h_{t-1}+\bar{B}_{t}x_{t},\quad y_{t}=C_{t}h_{t},$ (6)
@@
$\bar{A}=\exp(\Delta A),\quad\bar{B}=(\Delta A)^{-1}(\exp(\Delta A)-I)\cdot\Delta B,$ (7)
@@
Cross-Modal Global-Context. While BiMamba [8] excels at capturing local dependencies, we employ a Transformer [39] block to model the cross-modal global context, which is crucial for aligning the overall dance structure with long-term musical phrasing. This block consists of a cross-attention layer followed by a feed-forward network (FFN). The attention layer is the core component as:
$\text{Attention}=\text{softmax}\left(\frac{Q_{d}\cdot K_{m}^{T}}{\sqrt{C}}\right)V_{m}.$ (8)
Table 1: Quantitative comparison with SOTAs on the MA-Data dataset in Music-Driven Dance Video Generation task.
@@
Hallo2 [6] [ICLR’25] 62.64 50.79 92.48 93.84 98.30 96.56 16.55 1.29 8.11 5.47 0.505
EDGE [17] [CVPR’23] 63.05 49.70 91.79 93.30 98.64 97.10 21.77 1.39 9.08 5.74 0.498
Lodge [19] [CVPR’24] 63.69 49.22 91.67 92.98 98.46 97.05 18.73 1.49 8.87 5.71 0.474
MEGA [47] [NeurIPS’25] 66.14 49.89 92.95 94.13 97.45 96.32 18.98 1.65 8.78 5.59 0.513
@@
Table 2: Quantitative comparison with SOTAs on the FineDance dataset in Music-Driven 3D Dance Generation task.
@@
FACT [16] [ICCV’21] 113.38 97.05 0.284 3.36 6.37 0.183 29
MNET [15] [CVPR’22] 104.71 90.31 0.394 3.12 6.14 0.186 26
Bailando [30] [CVPR’22] 82.81 28.17 0.188 7.74 6.25 0.202 188
EDGE [37] [CVPR’23] 94.34 50.38 0.200 8.13 6.45 0.212 119
Lodge [19] [CVPR’24] 50.00 35.52 0.028 5.67 4.96 0.226 224
MEGA [47] [NeurIPS’25] 50.00 13.02 0.243 6.23 6.27 0.226 238
@@
3.3 Appearance Expert
Wan-Animate [5] has recently garnered substantial attention in both industry and academia. However, it is designed for general-purpose motion synthesis; direct transfer to dance video generation is suboptimal due to the domain gap and the richer spatiotemporal complexity of dance—namely intricate whole-body coordination and dynamic camera choreography. Accordingly, the Appearance Expert adopts a decoupled Kinematic–Aesthetic fine-tuning strategy to achieve high-fidelity appearance synthesis for dance videos.
Model Architecture. As illustrated in Fig. 2, the architecture of our Appearance Expert is built upon the Wan-Animate [5], which takes a reference image $I$ for appearance and a 3D motion sequence $X$ for motion guidance. The motion sequence $X$ is first projected to 2D keypoints, which are then encoded by a Body Adapter to yield motion features. These features are subsequently fused with the latent extracted from the reference image $I$. The resulting latent is processed by a backbone of stacked DiT blocks, where lightweight LoRA adapters are integrated into each block. The facial processing pipeline remains identical to that of Wan-Animate and is therefore omitted for clarity.
Kinematic Stage. In dance, body pose is paramount. The original Wan‑Animate prioritizes facial cues, allocating a dedicated cross‑attention branch to the face while fusing body signals only via additive injection. We therefore strengthen kinematic conditioning by fine‑tuning the Body Adapter in the Kinematic Stage to reweight and calibrate body features across scales, thereby enforcing motion adherence without altering the backbone. We intentionally do not introduce an additional body cross‑attention branch because (i) it disturbs the pretrained inductive bias and can compete with the facial cross‑attention, causing feature entanglement, (ii) it adds substantial memory/latency overhead and training instability on long, fast dance sequences.
Aesthetic Stage. To refine visual quality without disturbing motion control, we freeze the kinematic pathways and attach lightweight LoRA adapters to the attention (query/key/value/output) and feed-forward projections in each DiT block of Wan‑Animate. These rank‑r adapters enable parameter‑efficient specialization toward dance‑specific aesthetics—sharpening textures (skin, hair, fabric), stabilizing clothing and accessories, and handling rich camera choreography (pans, zooms, handheld motion)—while preserving pretrained content priors. Specifically, LoRA is an effective technique for adapting large pre-trained models to down-streaming tasks with few training-able parameters. To achieve this goal, LoRA introduces a low-rank decomposition-based method to the model’s weight matrix, enabling efficient adaptation to new tasks while maintaining the model’s original capabilities. Given the weight matrix $W_{0}\in\mathbb{R}^{m\times n}$ of the original pre-trained model, LoRA [11] uses two low-rank matrices $A\in\mathbb{R}^{m\times r}$ ($r\ll m$) and $B\in\mathbb{R}^{r\times n}$ ($r\ll n$) to shift the trained distribution according to the new data training. Thanks to the low-rank matrix in $A$ and $B$, LoRA updates the model more efficiently than the full rank matrix and shows comparable results with full training. Formally, the new weight matrix $W$ can be represented as:
$W=W_{0}+\Delta W=W_{0}+AB.$ (9)
4 Experiment
4.1 Dataset
@@
To support music-driven dance video generation task, we curate a large-scale dance video dataset, named MA-Data. It comprises  70k clips of 5–10 seconds each (totaling 116 hours), and spans over 20 distinct dance genres, such as Jazz, Latin, Eastern Folk. MA-Data consists of two complementary sources. (1) 3D-rendered data (motion-centric). This subset is derived from FineDance [17], the largest 3D dance dataset recorded by professional dancers, and emphasizes dance motion rather than visual appearance. Specifically, we obtain front-view videos from 3D motion by 3D rendering technique, and extract random 5–10 s segments via a sliding window for data augmentation, yielding  20k clips ( 28 h). (2) In-the-wild internet data (appearance-centric). This subset is curated from high-engagement creators on platforms such as TikTok and YouTube, emphasizing visual appearance; motions typically prioritize entertainment value over technical rigor. As raw crawls include many samples misaligned with our task, we apply a multi-stage cleaning pipeline: (i) perform shot boundary detection with TransNet V2 [34], segment accordingly, and discard segments shorter than 5 s; (ii) remove near-static clips using an optical-flow magnitude threshold; (iii) enforce a single-performer constraint via ViTPose [41] by dropping clips containing two or more people; and (iv) split long videos into 5–10 s clips with a sliding window and random offsets. The final set comprises  50k clips ( 88 hours).

### v3

@@ 3.2. Motion Expert
DDPM (Ho et al., 2020) defines diffusion as a Markov noising process with latents $\{z_{t}\}_{t=0}^{T}$ that follow a forward noising process $q(z_{t}|x)$, where $x\sim p(x)$ is drawn from the 3D dance data distribution. The forward noising process is defined as:
(1) $q(z_{t}|x)\sim\mathcal{N}(\sqrt{\bar{\alpha}_{t}}x,(1-\bar{\alpha}_{t})I),$
@@
where $\bar{\alpha}_{t}\in(0,1)$ are constants which follow a monotonically decreasing schedule such that when $\bar{\alpha}_{t}$ approaches 0. Timestep $T$ are commonly set to 1000, and $z_{T}\sim\mathcal{N}(0,I)$. With paired music conditioning $c$, we can reverse the forward diffusion process by learning to estimate $\hat{x}_{\theta}(z_{t},t,c)\approx x$ with model parameters $\theta$ for all $t$. We can optimize $\theta$ by the naive reconstruction loss in Diffusion Model (Ho et al., 2020):
(2) $\mathcal{L}_{\text{DM}}=\mathbb{E}\big[\|\hat{x}_{\theta}(z_{t},t,c)-x\|_{2}^{2}\big].$
Guidance-Free Training.
Conventional classifier-free guidance (CFG (Ho and Salimans, 2022)) modifies the sampling distribution only at inference time by combining conditional and unconditional predictions, which can introduce distribution mismatch and insufficient optimization toward the guided target distribution. In contrast, Guidance-Free Training (GFT (Chen et al., 2025c)) retains the same maximum-likelihood training objective as CFG but adopts a different parameterization that enables a single model to implicitly represent temperature-controlled sampling behavior during training, thereby mitigating distribution mismatch and yielding more stable and consistent high-fidelity generation. Accordingly, we establish $x_{\beta}$ as the new optimization target for our model $\theta$:
(3) $x_{\beta}=\beta\hat{x}_{\theta}(z_{t},t,c,\beta)+(1-\beta)\mathbf{sg}[\hat{x}_{\theta}(z_{t},t,\emptyset,1)]$
@@
(4) $\mathcal{L}_{\text{rec}}$ $=\mathbb{E}\big[\|x_{\beta}-x\|_{2}^{2}\big],$
@@
(5) $\mathcal{L}=\lambda_{\text{rec}}\mathcal{L}_{\text{rec}}+\lambda_{\text{joint}}\mathcal{L}_{\text{joint}}+\lambda_{\text{vel}}\mathcal{L}_{\text{vel}}+\lambda_{\text{foot}}\mathcal{L}_{\text{foot}}.$
Inference.
@@
At each of the denoising timesteps $t$, Motion Expert predicts the denoised sample and noises it back to timestep t1t-1: $\hat{z}_{t-1}\sim q(\hat{x}_{\theta}(z_{t},t,c,\beta),t-1)$, terminating when it reaches $t=0$. We utilize Denoising Diffusion Implicit Models (DDIM (Song et al., 2021)) to accelerate the sampling procedure. Values of $\beta$ near 0 favor high fidelity, while values near 1 favor high diversity. Thus, $\beta$ can also be regarded as a control signal, and we set its value to 0.75. Notably, GFT inherently achieves theoretically double the generation efficiency compared to conventional CFG, as it only requires a single conditional computation per step, eliminating the need for simultaneous conditional and unconditional predictions.
3.2.2. Model Architecture
Overview.
@@ 3.2.2. Model Architecture
Motion Expert adopts a BiMamba–Transformer hybrid backbone, thereby enabling the generation of temporally coherent and musically aligned dance motions. BiMamba captures intra-modal local dependencies in music or dance, while the Transformer models cross-modal global context. As shown in Fig.  2, the architecture details are as follows: Firstly, our model conditions the generator on the Librosa (McFee et al., 2015)-extracted music features from $M$ as (Li et al., 2021), which are then processed by an $L_{m}$‑layer BiMamba to capture intra‑modal temporal dynamics. Secondly, the diffusion time step $t$ and temperature parameter $\beta$ are encoded as sinusoidal embeddings and fused by element-wise addition to yield a $t$-$\beta$ embedding used throughout the generator. Third, the dance generator consists of $L_{d}$ stacked blocks. In each block: (1) the current state $z_{t}$ is first passed through a BiMamba to model intra-modal local dependencies; (2) FiLM  (Perez et al., 2018) is applied to modulate the features with the fused $t$-$\beta$ embedding; (3) a Transformer performs cross-modal attention over the music encoding to integrate global musical context, and subsequently passes the result through a feed-forward network; and (4) a second FiLM (Perez et al., 2018) further reinforces the $t$-$\beta$ conditioning. Finally, the generator outputs the 3D motion sequence $\hat{x}_{\theta}(z_{t},t,c,\beta)$ (i.e. $X$ in Sec. 3.1 Overview), represented as SMPL (Loper et al., 2023) parameters. Owing to this architecture, the Motion Expert generates the entire sequence in a non-autoregressive manner during inference, not only improving generation efficiency but also avoiding the exposure-bias problem in autoregressive (Yang et al., 2025b) and inpainting-based (Tseng et al., 2023) methods.
Intra-Modal Local-Dependency.
@@
While the Transformer excels at temporal modeling, it is inherently position-invariant and captures sequence order only through positional encodings (Vaswani, 2017), which limits its deep understanding of local dependencies. In contrast, music-to-dance generation demands strong local continuity between movements. Owing to its inherent sequential inductive bias, Mamba (Gu and Dao, 2023) has demonstrated strong performance in modeling fine-grained local dependencies  (Xu et al., 2024b; Fu et al., 2024). Moreover, Bidirectional Mamba processes inputs in both forward and backward directions, enabling wider representations and deeper understanding of music and dance. Specifically, the Selective State Space Model (Mamba) integrates a selection mechanism and a scan module (S6) (Gu and Dao, 2023) to dynamically emphasize salient input segments for efficient sequence modeling. Unlike traditional SSMs with time-invariant parameters, Mamba generates input-dependent $\bar{A}_{t},\bar{B}_{t},C_{t}$ through fully connected layers, enhancing generalization. For each time step $t$, the input $x_{t}$, hidden state $h_{t}$, and output $y_{t}$ evolve as:
(6) $h_{t}=\bar{A}_{t}h_{t-1}+\bar{B}_{t}x_{t},\quad y_{t}=C_{t}h_{t},$
@@
(7) $\bar{A}=\exp(\Delta A),\quad\bar{B}=(\Delta A)^{-1}(\exp(\Delta A)-I)\cdot\Delta B,$
@@
Cross-Modal Global-Context.
While BiMamba (Gu and Dao, 2023) excels at capturing local dependencies, it is less effective at modeling cross-modal global interactions. Thus, we employ a Transformer (Vaswani, 2017) module after BiMamba in each denoising block, which is crucial for aligning the overall dance structure with long-term musical phrasing. This block consists of a cross-attention layer followed by a feed-forward network (FFN). In the cross-attention layer, motion features serve as queries, while music features provide keys and values:
(8) $\text{Attention}=\text{softmax}\left(\frac{Q_{d}\cdot K_{m}^{T}}{\sqrt{C}}\right)V_{m}.$
In this way, the two components play complementary roles: BiMamba stabilizes short-range intra-modal dynamics, while the Transformer injects cross-modal global musical context to align the generated motion with the overall rhythm and phrase structure.
Figure 3. Qualitative comparison with SOTAs across reference image domains (real-person vs. anime-character) and music genres (Eastern Folk vs. Popping) in the music-driven dance video generation task.
3.3. Appearance Expert
Wan-Animate (Cheng et al., 2025) has recently garnered substantial attention in both industry and academia. However, it is designed for general-purpose motion synthesis; direct transfer to dance video generation is suboptimal due to the domain gap and the richer spatiotemporal complexity of dance—namely intricate whole-body coordination and dynamic camera choreography. Accordingly, the Appearance Expert adopts a decoupled Kinematic–Aesthetic fine-tuning strategy to achieve high-fidelity appearance synthesis for dance videos. Specifically, the Kinematic Stage fine-tunes only the Body Adapter while freezing the remaining components, whereas the Aesthetic Stage fine-tunes only the LoRA parameters while keeping the rest of the network fixed.
Model Architecture.
As illustrated in Fig.  2, the architecture of our Appearance Expert is built upon the Wan-Animate (Cheng et al., 2025), which takes a reference image $I$ for appearance and a 3D motion sequence $X$ for motion guidance. The motion sequence $X$ is first projected to 2D keypoints, which are then encoded by a Body Adapter to yield motion features. These features are subsequently fused with the latent extracted from the reference image $I$. The resulting latent is processed by a backbone of stacked DiT blocks, where lightweight LoRA adapters are integrated into each block. The facial processing pipeline remains identical to that of Wan-Animate and is therefore omitted for clarity.
Projector.
We introduce a 3D-to-2D Motion Projector to convert the SMPL sequence generated by the Motion Expert into the 2D pose format required by Wan-Animate. For each frame, we first transform the SMPL parameters into a 3D mesh and render it with pyrender under a fixed frontal-view camera, then apply ViTPose (Xu et al., 2022) to extract the corresponding 2D keypoints. In this way, the projector preserves the benefits of 3D motion modeling while enabling seamless integration with the downstream Appearance Expert.
Kinematic Stage.
In dance, body pose is paramount. The original Wan‑Animate prioritizes facial cues, allocating a dedicated cross‑attention branch to the face while fusing body signals only via additive injection. We therefore strengthen kinematic conditioning by fine‑tuning the Body Adapter in the Kinematic Stage to reweight and calibrate body features across scales, thereby enforcing motion adherence without altering the backbone. We intentionally do not introduce an additional body cross‑attention branch because (i) it disturbs the pretrained inductive bias and can compete with the facial cross‑attention, causing feature entanglement, (ii) it adds substantial memory/latency overhead and training instability on long, fast dance sequences.
Aesthetic Stage.
To refine visual quality without disturbing motion control, we freeze the kinematic pathways and attach lightweight LoRA adapters to the attention (query/key/value/output) and feed-forward projections in each DiT block of Wan‑Animate. These rank‑r adapters enable parameter‑efficient specialization toward dance‑specific aesthetics—sharpening textures (skin, hair, fabric), stabilizing clothing and accessories, and handling rich camera choreography (pans, zooms, handheld motion)—while preserving pretrained content priors. Specifically, LoRA is an effective technique for adapting large pre-trained models to down-streaming tasks with few training-able parameters. To achieve this goal, LoRA introduces a low-rank decomposition-based method to the model’s weight matrix, enabling efficient adaptation to new tasks while maintaining the model’s original capabilities. Given the weight matrix $W_{0}\in\mathbb{R}^{m\times n}$ of the original pre-trained model, LoRA (Hu et al., 2022) uses two low-rank matrices $A\in\mathbb{R}^{m\times r}$ ($r\ll m$) and $B\in\mathbb{R}^{r\times n}$ ($r\ll n$) to shift the trained distribution according to the new data training. Thanks to the low-rank matrix in $A$ and $B$, LoRA updates the model more efficiently than the full rank matrix and shows comparable results with full training. Formally, the new weight matrix $W$ can be represented as:
(9) $W=W_{0}+\Delta W=W_{0}+AB.$
Table 1. Quantitative comparison with SOTAs on the MA-Data dataset in Music-Driven Dance Video Generation task.
@@ 3.3. Appearance Expert
Hallo2 (Cui et al., 2024) [ICLR’25] 62.64 50.79 92.48 93.84 98.30 96.56 16.55 1.29 8.11 5.47 0.505
WAN-S2V (Gao et al., 2025) [Arxiv’25] 64.10 50.20 92.30 93.40 98.20 96.70 18.90 1.45 7.60 5.44 0.485
Echomimic-V3 (Meng et al., 2025) [AAAI’26] 63.20 49.00 91.90 93.10 98.05 96.40 19.60 1.32 7.20 4.60 0.460
EDGE (Li et al., 2023) [CVPR’23] 63.05 49.70 91.79 93.30 98.64 97.10 21.77 1.39 9.08 5.74 0.498
Lodge (Li et al., 2024c) [CVPR’24] 63.69 49.22 91.67 92.98 98.46 97.05 18.73 1.49 8.87 5.71 0.474
MEGA (Yang et al., 2025b) [NeurIPS’25] 66.14 49.89 92.95 94.13 97.45 96.32 18.98 1.65 8.78 5.59 0.513
@@
Table 2. Quantitative comparison with SOTAs on the FineDance dataset in Music-Driven 3D Dance Generation task.
@@
FACT (Li et al., 2021) [ICCV’21] 113.38 97.05 0.284 3.36 6.37 0.183 29
MNET (Kim et al., 2022) [CVPR’22] 104.71 90.31 0.394 3.12 6.14 0.186 26
Bailando (Siyao et al., 2022) [CVPR’22] 82.81 28.17 0.188 7.74 6.25 0.202 188
EDGE (Tseng et al., 2023) [CVPR’23] 94.34 50.38 0.200 8.13 6.45 0.212 119
Lodge (Li et al., 2024c) [CVPR’24] 50.00 35.52 0.028 5.67 4.96 0.226 224
MEGA (Yang et al., 2025b) [NeurIPS’25] 50.00 13.02 0.243 6.23 6.27 0.226 238
@@
4. Experiment
4.1. Dataset
@@
To support music-driven dance video generation task, we curate a large-scale dance video dataset, named MA-Data. It comprises  70k clips of 5–10 seconds each (totaling 116 hours), and spans over 20 distinct dance genres, such as Jazz, Latin, Eastern Folk. MA-Data consists of two complementary sources. (1) 3D-rendered data (motion-centric). This subset is derived from FineDance (Li et al., 2023), the largest 3D dance dataset recorded by professional dancers, and emphasizes professional dance motion rather than visual appearance. Specifically, we first retarget the motion sequence to a character model, then render front-view videos from the 3D character, and extract random 5–10 s segments with a sliding-window strategy for data augmentation, yielding  20k clips ( 28 hours). (2) In-the-wild internet data (appearance-centric). This subset is curated from high-engagement creators on platforms such as TikTok and YouTube, emphasizing visual appearance; motions typically prioritize entertainment value over technical rigor. As raw crawls include many samples misaligned with our task, we apply a multi-stage cleaning pipeline: (i) perform shot boundary detection with TransNet V2 (Soucek and Lokoc, 2024), segment accordingly, and discard segments shorter than 5 s; (ii) remove near-static clips using an optical-flow magnitude threshold; (iii) enforce a single-performer constraint via ViTPose (Xu et al., 2022) by discarding clips that contain multiple people or exhibit little to no human motion; and (iv) split long videos into 5–10 s clips with a sliding window and random offsets. The final set comprises  50k clips ( 88 hours). Finally, we collect an additional 200 5-second clips to construct the test set, with high engagement on TikTok and across multiple dance genres.
4.2. Evaluation
The key challenges of music-driven dance video generation are (Yang et al., 2024c; Zhang et al., 2025a): (1) generating dance motions that are kinematically plausible while artistically expressive; and (2) achieving high-fidelity visual appearance with strong spatiotemporal consistency. Inspired by this, we introduce a motion–appearance evaluation protocol. (1) Motion dimension. We extract 2D keypoint sequences using ViTPose (Xu et al., 2022) from the dance videos and evaluate from a Human-Kinematics perspective. To evaluate the fidelity and diversity, we report FID and DIV across two feature spaces (Li et al., 2021; Siyao et al., 2022): (1) kinetic (k), capturing motion dynamics, and (2) geometric (g), encoding spatial joint relations. To measure music–motion synchronization, we utilize the Beat Alignment Score (BAS) (Li et al., 2021, 2024c). (2) Appearance dimension. Inspired by  (Chen et al., 2025d; Li et al., 2025; Ling et al., 2025), we adopt VBench (Huang et al., 2024)—a widely used benchmark in video generation—and select a set of dance-specific metrics. Our evaluation includes imaging quality (IQ), aesthetic quality (AQ), subject consistency (SC), background consistency (BC), motion smoothness (MS), temporal flickering (TF).